# A kiloparsec-scale ordered magnetic field in a galaxy at z=5.6

### v2

@@ 4.3 Formation timescales of galactic magnetic fields
where is the volume density of the ISM, and is the velocity dispersion of the gas. In SPT0346-52, the observed [C ii] line dispersion at the region with polarisation detection is about 140 km s-1 (see also Fig. 5). We assume a preferred gas density of 50 cm-3 from the photodissociation regions (PDR) modelling (Litke et al. 2022). We estimate the equipartition turbulent magnetic field strength to be G. We emphasise that this is only a theoretical upper limit for the turbulent magnetic fields, with the assumption that the turbulence dominates the line width. If the line width is dominated by regular rotation or the velocity difference of the possible two merging galaxies, the equipartition field strength can be several times lower, which will bring the field strength closer to the local starburst galaxies, either measured with dust polarisation (Lopez-Rodriguez et al. 2021; Lopez-Rodriguez 2023) or radio synchrotron observations (Lacki & Beck 2013). It should be noted that the magnetic field strength in the cold phase of the interstellar medium is different from that of the warm and diffuse medium measured at radio wavelengths (Martin-Alvarez et al. 2023). The field strength in the cold and dense medium could be larger than that from the warm and diffuse medium, as the field strength correlates with the gas volume density. Nonetheless, our calculation supports that highly amplified magnetic fields could be developed in very young star-forming galaxies.
@@ 4.4 Implications for early galaxy evolution
Magnetic fields have been widely accepted as an essential part of galaxy evolution, but their importance is still under debate, especially on galactic scales. For instance, to affect the gas collapse/infall on a free-fall timescale, the ordered field strength needs to be G (e.g. Ceverino et al. 2012), which is 10100 larger than most observations. However, Krumholz & Federrath (2019) summarised the effects of magnetic fields in various star-formation processes, concluding that they are not the dominant player in controlling star formation inside galaxies, but they could play a larger indirect role by the interaction with stellar feedback and cosmic rays. Such indirect intervention will not affect galaxy evolution in a short timescale but could be substantial over long cosmic time. Our observations prove galactic magnetic fields are already present in the Cosmic Dawn, which supports the picture that galactic magnetic fields could co-evolve with the host galaxy over a considerable cosmic period.
@@
A quickly formed magnetic field could also intervene with the earliest galaxy formation. James Webb Space Telescope (JWST) has recently reported an overabundance of massive galaxies at (Harikane et al. 2023; Finkelstein et al. 2023; Labbé et al. 2023). The possible explanations proposed so far include Pop-III stars (e.g. Ventura et al. 2024), top-heavy initial mass functions (IMFs) (e.g. Yung et al. 2024), feedback-free starbursts (Dekel et al. 2023), early positive AGN feedback (e.g. Silk et al. 2024) or simply the incorrect modelling of their stellar masses (e.g. Wang et al. 2024). The quick formation of proto-galaxies also indicates that various dynamos could start much earlier and quicker, leading to an early formation of G magnetic fields in the ISM. We, therefore, propose that magnetic fields could be another viable factor to impact the formation of early massive galaxies, including the processes such as enhancing the formation of massive stars and modifying the feedback process of the starburst and AGN (e.g. Sharda et al. 2020; Sharda & Menon 2024; Stacy et al. 2022; Sadanari et al. 2024; Begelman & Silk 2023). In the future, surveys of magnetic fields in more distant galaxies could provide better statistical evidence about the importance of magnetic fields in the early Universe.
@@
Meanwhile, the rapid formation of galactic magnetic fields could also play an important role on a larger scales. A statistical study of nearly 200 nearby galaxies has found that magnetic field strength decreases as a function of distance from the galactic disk, but it is enhanced in the bipolar direction, indicating that the CGM is magnetised by the galactic outflow (Heesen et al. 2023). With sensitive low-frequency telescopes, like Low Frequency Array (LOFAR), magnetic fields have been observed up to Mpc around the galaxy cluster (Di Gennaro et al. 2020; Cuciti et al. 2022). The early development of magnetic fields in the ISM supports the hypothesis that the magnetic fields could be transferred from the ISM to the CGM (Bertone et al. 2006; Hanasz et al. 2013). This process could be rather efficient, along with the ubiquitous presence of galactic outflows found in dusty star-forming galaxies (e.g. Spilker et al. 2020), which can provide a natural explanation for large-scale scale-magnetic fields in the CGM of local galaxies. polarised dust emission extends over 3 kpc and shows a per cent variance. We measure a bimodal distribution of the magnetic field position angle, peaking at 28 degrees and 78 degrees. The region with polarisation detection is spatially coincident with the [C ii] emission around a velocity component of km s-1, and it also has a generally small velocity gradient and velocity dispersion. Our observations demonstrate that mapping and resolving kpc-scale polarised dust emission back to the Cosmic Dawn is possible with modern submillimetre interferometric observations.
@@
The observed kpc-scale polarised dust emission indicates the presence of large-scale ordered magnetic fields. We discuss the formation and implications of such large-scale ordered magnetic fields. Combining all the available observations, the general properties of SPT034652, and the field structures in the source plane favour a galaxy merger event as the primary driver of the kpc-scale ordered magnetic fields. With the assumption that the observed ordered magnetic fields are stretched from the existing magnetic fields of the pre-merged galaxies, our observation supports kpc-scale ordered magnetic fields can form in the early massive star-forming galaxy by the time the Universe is only 1 Gyr old. Assuming equipartition with the turbulent kinetic energy, we also compute the upper limit of the turbulent magnetic field strength of G, which could be amplified by the extreme star formation activity. The quick formation of magnetic fields provides new implications for forming massive galaxies in the very early Universe. Such early magnetic fields could mediate the early star formation and play a crucial role in the long-term evolution of galaxies. The early presence of magnetic fields in the interstellar medium could also contribute to the formation of large-scale magnetic fields in the circum-galactic medium.
@@
The future ahead is promising. The current results are biased to a single extreme starburst galaxy, but this approach can be easily applied to many other strongly lensed dusty star-forming galaxies (DSFGs) and even the brightest unlensed targets in the early Universe with a reasonable amount of telescope time. Future surveys on a large sample of DSFGs at different redshifts and star formation rates could be used to test whether the magnetic fields are common in the ISM of young star-forming galaxies and how they evolve across cosmic time. In addition, high-resolution observations of the best available targets will reveal more detailed structures on smaller scales, which will be crucial to understanding the possible mechanisms driving their early development.
@@ Acknowledgements.
We thank the anonymous referee for providing useful comments, which significantly improved this manuscript. J.C. is grateful for the useful help from Andy Biggs and Johan Richard, which has inspired this project. J.C. gratefully acknowledges the support provided by European Research Council, grant number 101055023. E.L.-R. is supported by the NASA/DLR Stratospheric Observatory for Infrared Astronomy (SOFIA) under the 08_0012 Program. SOFIA is jointly operated by the Universities Space Research Association,Inc.(USRA), under NASA contract NNA17BF53C, and the Deutsches SOFIA Institut (DSI) under DLR contract 50OK0901 to the University of Stuttgart. E.L.-R. is supported by the NASA Astrophysics Decadal Survey Precursor Science (ADSPS) Program (NNH22ZDA001N-ADSPS) with ID 22-ADSPS22-0009 and agreement number 80NSSC23K1585.This paper makes use of the following ALMA data: 2021.1.00458.S and 2013.1.01231.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile.
[…]

### v3

@@ 4.3 Formation timescales of galactic magnetic fields
where is the volume density of the ISM and is the velocity dispersion of the gas. In SPT0346-52, the observed [C ii] line dispersion at the region with polarisation detection is about 140 km s-1 (see also Fig. 5). We assume a preferred gas density of 50 cm-3 from photodissociation region modelling (Litke et al. 2022). We estimate the equipartition turbulent magnetic field strength to be G. We emphasise that this is only a theoretical upper limit for the turbulent magnetic fields, with the assumption that the turbulence dominates the line width. If the line width is dominated by regular rotation or the velocity difference of the two possible merging galaxies, the equipartition field strength can be several times lower, which will bring the field strength closer to the local starburst galaxies, measured with either dust polarisation (Lopez-Rodriguez et al. 2021; Lopez-Rodriguez 2023) or radio synchrotron observations (Lacki & Beck 2013). It should be noted that the magnetic field strength in the cold phase of the ISM is different from that of the warm and diffuse medium measured at radio wavelengths (Martin-Alvarez et al. 2023). The field strength in the cold and dense medium could be larger than that from the warm and diffuse medium, as the field strength correlates with the gas volume density. Nonetheless, our calculation supports the idea that highly amplified magnetic fields could be developed in very young SFGs.
@@ 4.4 Implications for early galaxy evolution
Magnetic fields have been widely accepted as an essential part of galaxy evolution, but their importance is still under debate, especially on galactic scales. For instance, to affect the gas collapse or infall on a free-fall timescale, the ordered field strength needs to be G (e.g. Ceverino et al. 2012), which is 10100 larger than most observations. However, Krumholz & Federrath (2019) summarised the effects of magnetic fields in various star-formation processes, concluding that they are not the dominant player in controlling star formation inside galaxies, but that they could play a larger indirect role by interacting with stellar feedback and cosmic rays. Such an indirect intervention will not affect galaxy evolution over a short timescale but could be substantial over long cosmic time. Our observations prove that galactic magnetic fields were already present at the Cosmic Dawn, which supports the picture that galactic magnetic fields could co-evolve with the host galaxy over a considerable cosmic period.
@@
A quickly formed magnetic field could also intervene with the earliest galaxy formation. The James Webb Space Telescope (JWST) has recently reported an overabundance of massive galaxies at (Harikane et al. 2023; Finkelstein et al. 2023; Labbé et al. 2023). The possible explanations proposed so far include Population III stars (e.g. Ventura et al. 2024), top-heavy initial mass functions (e.g. Yung et al. 2024), feedback-free starbursts (Dekel et al. 2023), early positive AGN feedback (e.g. Silk et al. 2024), or simply incorrect modelling of their stellar masses (e.g. Wang et al. 2024). The quick formation of proto-galaxies also indicates that various dynamos could start much earlier and quicker, leading to the early formation of G magnetic fields in the ISM. We therefore propose that magnetic fields could be another viable factor to impact the formation of early massive galaxies, including processes such as enhancing the formation of massive stars and modifying the feedback process of the starburst and AGN (e.g. Sharda et al. 2020; Sharda & Menon 2024; Stacy et al. 2022; Sadanari et al. 2024; Begelman & Silk 2023). In the future, surveys of magnetic fields in more distant galaxies could provide better statistical evidence about the importance of magnetic fields in the early Universe.
@@
Meanwhile, the rapid formation of galactic magnetic fields could also play an important role on larger scales. A statistical study of nearly 200 nearby galaxies has found that magnetic field strength decreases as a function of distance from the galactic disk, but is enhanced in the bipolar direction, indicating that the CGM is magnetised by the galactic outflow (Heesen et al. 2023). With sensitive low-frequency telescopes, like the Low Frequency Array (LOFAR), magnetic fields have been observed up to Mpc around the galaxy cluster (Di Gennaro et al. 2020; Cuciti et al. 2022). The early development of magnetic fields in the ISM supports the hypothesis that magnetic fields could be transferred from the ISM to the CGM (Bertone et al. 2006; Hanasz et al. 2013). This process could be rather efficient, along with the ubiquitous presence of galactic outflows found in dusty SFGs (e.g. Spilker et al. 2020), which can provide a natural explanation for large-scale magnetic fields in the CGM of local galaxies.
5 Conclusions
In this paper, we report the detection of linearly polarised thermal dust emission in SPT034652 at . The polarised dust emission extends over 3 kpc and shows a per cent variance. We measure a bimodal distribution of the magnetic field position angle, peaking at 28 degrees and 78 degrees. The region with polarisation detection is spatially coincident with the [C ii] emission around a velocity component of km s-1, and it also has a generally small velocity gradient and velocity dispersion. Our observations demonstrate that mapping and resolving kpc-scale polarised dust emission back to the Cosmic Dawn is possible with modern submillimetre interferometric observations.
@@ 5 Conclusions
The observed kpc-scale polarised dust emission indicates the presence of large-scale ordered magnetic fields. We discuss the formation and implications of such large-scale ordered magnetic fields. Combining all the available observations, the general properties of SPT034652 and the field structures in the source plane favour a galaxy merger event as the primary driver of the kpc-scale ordered magnetic fields. With the assumption that the observed ordered magnetic fields are stretched from the existing magnetic fields of the pre-merged galaxies, our observation supports kiloparsec-scale ordered magnetic fields being able to form in the early massive star-forming galaxy by the time the Universe is only 1 Gyr old. Assuming equipartition with the turbulent kinetic energy, we also computed the upper limit of the turbulent magnetic field strength of G, which could be amplified by the extreme star formation activity. The quick formation of magnetic fields provides new implications for forming massive galaxies in the very early Universe as they could mediate the early star formation and play a crucial role in the long-term evolution of galaxies. Furthermore, the early presence of magnetic fields in the interstellar medium could also contribute to the formation of large-scale magnetic fields in the circum-galactic medium.
@@
The future is promising. The current results are biased towards a single extreme starburst galaxy, but this approach could easily be applied to many other strongly lensed dusty star-forming galaxies and even the brightest unlensed targets in the early Universe with a reasonable amount of telescope time. Future surveys of a large sample of DSFGs at different redshifts and star formation rates could be used to test whether the magnetic fields are common in the ISM of young star-forming galaxies and how they evolve across cosmic time. In addition, high-resolution observations of the best available targets will reveal more detailed structures on smaller scales, which will be crucial to understanding the possible mechanisms driving their early development.
@@ Acknowledgements.
We thank the anonymous referee for providing useful comments, which significantly improved this manuscript. J.C. is grateful for the useful help from Andy Biggs and Johan Richard, which has inspired this project. J.C. gratefully acknowledges the support provided by European Research Council, grant number 101055023. E.L.-R. is supported by the NASA/DLR Stratospheric Observatory for Infrared Astronomy (SOFIA) under the 08_0012 Program. SOFIA is jointly operated by the Universities Space Research Association,Inc.(USRA), under NASA contract NNA17BF53C, and the Deutsches SOFIA Institut (DSI) under DLR contract 50OK0901 to the University of Stuttgart. E.L.-R. is supported by the NASA Astrophysics Decadal Survey Precursor Science (ADSPS) Program (NNH22ZDA001N-ADSPS) with ID 22-ADSPS22-0009 and agreement number 80NSSC23K1585. SD is supported by the UK STFC consolidated grant ST/X000982/1.This paper makes use of the following ALMA data: 2021.1.00458.S and 2013.1.01231.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile.
This research made use of Astropy, a community-developed core Python package for Astronomy (Astropy Collaboration et al. 2022); NumPy, a fundamental package for scientific computing with Python; Matplotlib, a plotting library for Python; IPython, an interactive computing system for Python; photutils, an affiliated package of AstroPy to provide tools for detecting and performing photometry of astronomical sources (Bradley et al. 2020).